OpenStudy (anonymous):

Algebra: PLEASE HELP !? 1) Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. A) –2x – y = –7 B) 2x – y = –7 C) 2x – y = 11 D) –2x + y = –11 2) Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. A) y – 4 = 2(x + 3) B) y + 4 = 2(x – 3) C) y + 3 = 2(x – 4) D) y – 3 = 2(x + 4) 3) Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. A) 4x + 4y = 6 B) x + 4y = 12 C) 4x + 4y = –12 D) x – 4y = –6

OpenStudy (anonymous):

4) Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. A) y = 3x – 6 B) y = 3x – 3 C) y = 3x + 3 D) y = 3x + 6 I am so bad at this . Please help me!!

OpenStudy (anonymous):

to many and to lazy lol

OpenStudy (anonymous):

y=mx+B

OpenStudy (anonymous):

what is that ?

OpenStudy (anonymous):

the formula for finding slope of a line.

OpenStudy (anonymous):

what does it all stand for ?

OpenStudy (anonymous):

slope is equal to x

OpenStudy (anonymous):

you're going to have to walk me through this, it's like trying to read jibberish. i don't get this stuff at all !

OpenStudy (anonymous):

IT IS THE FORMULA FOR SLOPE WHERE M IS THE SLOPE X AND Y ARE THE POINTS ON THE GRAPH B IS THE Y-INTERCEPT there you go

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20080813164738AAv8I3w maybe this will help?

OpenStudy (anonymous):

idk what any of those words mean but i'll go to the link. thanks.

OpenStudy (anonymous):

did you not learn any of that and your teacher assigned this to you for homework? hopefully the link will help.

OpenStudy (anonymous):

i didn't get it & i'm homeschooled so i couldn't ask cos it's too late at night & she stops answering q's at 4pm

OpenStudy (anonymous):

oh okay, well if you want to be able to successfully solve those types of problems, it is necessary to know about y=mx+B. so basically just plug in the points into the equation(label them as x and y) and you'll find the answer. sorry if I'm not being very specific, let me know if you need it clarified any more.

OpenStudy (anonymous):

thank you very much !

OpenStudy (anonymous):

no problem!